I am confused as to what system to use on this problem When an owner spent $2000/month on ads and the revenue was 120,000 How much do the owner spend if the revenue was 180,000.00
Answer by stanbon(75887)

When an owner spent $2000/month on ads and the revenue was 120,000 How much do the owner spend if the revenue was 180,000.00
-----------------
x/180,000 = 2000/120,000
x=180,000(1/60)
x=3000
He would spend $3000 on ads.
Cheers,
Stan H.